MY COMPANY I CHOSE IS APPLE..... I ONLY NEED THE ANSWER FOR LIQUIDITY- CURRENT AND QUICK RATIO AND ASSET MANAGEMENT- TOTAL ASSETS TURNOVER AND FIXED ASSETS TURNOVER

As the new financial manager of your company, the CEO has asked your team to provide a brief

analysis of the company’s performance to present at the upcoming board of directors meeting.

The CEO has asked that you assess the company’s performance against your company’s

industry. Thus, to do this, you will need to use ratio analysis or other techniques to determine

areas in which the company is doing well, as well as areas that management should look at.

Here are the steps for the project:

1. Select your teammate. Each team should be made up of two or three members.

2. Determine which company you will analyze for the project. Your selection may be

subject to your professor’s approval. The company that you select will likely be used

for all four team projects. As such, be sure that the company has debt on its balance

sheet, as this will be a requirement for future projects.

3. Go to the website for your company and download the 10-K report for the most

recent year.

4. Perform your ratio analysis on your company:

a. A good place to start would be to perform a complete DuPont analysis of the

company. The DuPont analysis might provide guidance as to what particular

areas of the company should be examined next and what ratios should be

calculated. Be sure to include ratios that cover the following areas:

i. Profitability

ii. Debt Management

iii. Liquidity

iv. Asset Management

v. Market Value

b. In addition to the DuPont analysis ratios, be sure to present and discuss at

least six relevant ratios that your team feels may best assess the company’s

performance.

c. Prepare the same ratios that your team prepared above, but this time for the

prior year.

d. Provide an analysis that compares your company’s current year ratios to its

prior year ratios. There is no need to explain the purpose of the ratios
Answer #1

The latest 10-k report that is available for Apple is for the year 2018 (i.e. year ended September 29, 2018)

The ratios and its computations are shown below:

(1) Current ratio = current assets/current liabilities

A Total current assets 131,339.00
B Total current liabilities 116,866.00
C Current ratio = A/B 1.12

(2) Quick ratio = (Total current assets - inventory - prepaid expenses)/current liabilities

A Total current assets 131,339.00
B Inventories 3,956.00
C Total current liabilities 116,866.00
D Quick ratio = (A-B)/C 1.09

(3): Total asset turnover = Net sales/average total assets

Assets for 2018 Assets for 2017
365,725.00 375,319.00
A Average total assets 370,522.00
B Net sales 265,595.00
Asset turnover = B/A 0.72

(4): Fixed assets turnover = Net sales/Average net fixed assets

Fixed assets for 2018 Fixed assets for 2017
234,386.00 246,674.00
A Average fixed assets (net) 240,530.00
B Net sales 265,595.00
Fixed assets turnover = B/A 1.10